OpenStudy (anonymous):

For the function f(x) = 2 – 3x, find f(–4). A. –5 B. 14 C. –10 D. 12

OpenStudy (anonymous):

plug in -4 onto x and solve

OpenStudy (anonymous):

f(x) = 2 – 3(-4) = ?

OpenStudy (anonymous):

i dont get it..

OpenStudy (anonymous):

f(x) = 2 – 3(-4) = 2--12 = 2 + 12 = ?

OpenStudy (anonymous):

so 14?

OpenStudy (anonymous):

yea